OpenStudy (anonymous):

Please help me. Find the # of decibels for the power of the sound.Round to the nearest decibel. A rock concert, 5.37 x 10 to the negative power watts/cm square. D=_____dB

OpenStudy (shane_b):

It looks some info is missing:\[\large 5.37x10^{-?}\]

OpenStudy (anonymous):

-6...I'm sorry. Thanks

OpenStudy (shane_b):

Use the formula: \[dB=10log_{10}(\frac{I}{I_0})=10log_{10}(\frac{5.37x10^{-2}W/m^2}{10^{-12}})=?\] Note that I converted W/cm^2 to W/m^2.

OpenStudy (shane_b):

That should be 1x10^12 on the bottom...and I got 107.23 dB

OpenStudy (anonymous):

Can you explain the -2 next to the ten on top of the fraction and the negative twelve next to the 10 on the bottom of the fraction.. thanks for your assistance !

OpenStudy (shane_b):

For the top: \[1 m^2 = (100cm)^2=10,000cm^2\]Therefore, you need to shift the decimal place to the left 4 places to convert cm^2 to m^2:\[5.37x10^{-6+4}=5.37x10^{-2}\] The bottom number is just a constant that's part of the formula:\[I_0=1x10^{-12} \]

OpenStudy (anonymous):

thanks so much i really appreciate your help

OpenStudy (shane_b):

my pleasure